The angle of incidence is 40Β°, what is the value of the angle of refraction when the light travels from water to light flint glass?

1
Expert's answer
2021-11-14T17:10:38-0500

The angle of refraction is


"\theta_2=\arcsin(\sin\theta_1·\frac{n_1}{n_2}),\\\space\\
\theta_2=\arcsin(\sin40°·\frac{1.33}{1.66})=31°"